OpenStudy (anonymous):

Point P is the center of two concentric circles. PQ = 4.2 and PS = 8.2. RS is tangent to the smaller circle and a chord of the larger circle. What is length of RS to the nearest tenth? 12.2 8.1 16.4 14.1

OpenStudy (anonymous):

nope

Directrix (directrix):

Segment PQ is perpendicular to segment RS because a radius drawn to a tangent to a circle is perpendicular to the tangent at the point of tangency. That creates a right triangle and thoughts of the Pythagorean Theorem come to mind.

OpenStudy (anonymous):

i got B

Directrix (directrix):

@ehmuleeee If you use the Pythagorean Theorem to find QS, then we have one more theorem to apply.

Directrix (directrix):

@evilmath Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

Directrix (directrix):

@ehmuleeee I'm asking you to solve this: (8.2) ^2 = (4.2)^2 + x^2 where x = QS.

OpenStudy (anonymous):

67.24=17.64+ x^2

Directrix (directrix):

Right. 67.24=17.64+ x^2 x^2 = 67.24 - 17. 64 x^2 = ? @ehmuleeee

OpenStudy (anonymous):

49.6

Directrix (directrix):

x^2 = 49.6 x = positive square root of 49.6 x = ? @ehmuleeee

OpenStudy (anonymous):

7.04

Directrix (directrix):

Yes. If a diameter is perpendicular to a chord, it bisects the chord and its arcs. Check out the attached diagram. So, RS = QS = 2* 7.04 = ? @ehmuleeee

OpenStudy (anonymous):

14.1!

Directrix (directrix):

Correct. Note that the correct answer is not the one @evilmath threw out and left without explanation.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

to both of you.

OpenStudy (anonymous):

i appreciate your time & effort

Directrix (directrix):

You are welcome.